# MonkeyTree: Near-Minimal Congestion for Multi-tenant Training via Migration

### v1

@@ 2.1. Multi-Tenant Training and Congestion
2.2. Migration Challenges in CPU Workloads
@@ 3. Tractability of GPU Migration
DP replicates the model across GPUs, with each GPU processing a different data batch. Fully Sharded Data Parallelism (FSDP) extends DP by sharding parameters across GPUs. Both require synchronizing gradients or parameters at the end of each training iteration through collective communication. These collectives generate traffic proportional to model size with the same volume and pattern, so we treat them identically and refer to both as DP traffic. When combined with other parallelisms such as PP and TP, only a subset of GPUs shares the same model replica; we call this subset a replica group, and each performs its own collective independently.
@@ 3.1. Congestion-Free States are Achievable
This sparsity resolves the ill-defined optimization problem from Section 2.2. In particular, there exist states in which the traffic leaving each ToR does not exceed its uplink capacity (bottom-left of Figure 3). We define such states as congestion-free, since this condition allows each DP flow to be assigned a dedicated uplink, achieving path isolation. Under minimal uplink provisioning, such states always exist (§5.1). For CPU workloads, operators must balance network contention against migration overhead (joint_migration_routing; joint_migration_routing_2; joint_placement_routing_3). In contrast, for ML workloads this tradeoff disappears: because congestion-free states are always attainable, they can be enforced as a constraint, allowing the optimization to focus solely on minimizing migrations.
@@ 3.2. Congestion-Free States Are Abundant
3.3. Congestion-Free States are Self-Reinforcing
@@ 5.2. MonkeyTree ILP for DP and FSDP Jobs
We attribute this efficiency to the ability of branch-and-bound solvers (branch_bound) to exploit the abundance of the solution space. To build intuition for the runtime, let $\gamma$ denote the number of migrations in searching for a solution. The number of distinct cluster states reachable within $\gamma$ moves is $O(N^{2\gamma})$, where $N$ is the cluster size. As branch-and-bound progressively discovers feasible solutions with increasing migration counts, large portions of the search space are pruned, since the solver discards solutions requiring the same or more moves. We believe that, because the optimal number of migrations is typically small (§6.5), the solver explores a limited effective search space, resulting in fast ILP solving time.
@@ 5.3. Extending to Model Parallel Jobs
Pipeline parallelism (PP).   Pipeline parallelism shards the model between layers into distinct stages that run on separate GPUs. Activations are communicated point-to-point between stages. Data-parallel replicas are thus across each stage, and the subsequent all-reduce (DP-flows) only happens within these stages. We only use our Perfect routing algorithm to load balance pipeline traffic and ignore accommodating them in our formulation, because they are small (§3) and therefore have little impact on end-to-end job completion time. Instead, we update our formulation to treat pipeline stages as independent jobs, thereby controlling each stage’s DP-flow separately.
@@ 6.5. Migration Solver Scalability
Figure 9 presents the distribution of required job migration (number of moves) to lower the fragmentation below two on clusters of varying sizes across five traces each. The distribution shifts as cluster size increases, but the average remains consistently between $1.7$ and $1.8$. Approximately $80\%$ of the weight is concentrated on one or two moves across all scenarios, and move counts greater than $5$ are rare, occurring in fewer than $1\%$ of the migrations for each cluster.
@@
Finally, Figure 12 plots the ILP’s solve time across cluster sizes ranging from $128$ to $2,048$ GPUs. For clusters of $1,024$ GPUs or fewer, the average solve time remains under $0.5$ s with a p90 below $1$ s. At $2,048$ GPUs, solve time increases by over $5\times$, with an average of $2.39$ s and a p90 of $4.04$ s. Even so, this remains well within practical limits: migrations are triggered infrequently and can tolerate seconds of planning latency, making the ILP viable even at the large cluster sizes.
@@ 6.7. Threshold Sensitivity
Figure 13 plots the number of migrations alongside the average, p90, and p99 slowdown. At a threshold of $4$, matching the number of uplinks, p99 slowdown is within 66% of ideal, and the average is within 11%. This is slightly worse than the results in §6.2, as a workload composed entirely of FSDP and DP jobs with few uplinks presents a more constrained setting than the full workload mix. Decreasing the threshold below four yields less than 0.50.5% additional improvement: Perfect routing already guarantees path isolation at a fragmentation threshold of four, so further migrations cannot eliminate elephant-flow collisions. A lower threshold only reduces collisions between elephant flows and in-flight migration traffic, yielding a marginal benefit offset at the cost of up to $2\times$ more migrations.
@@ 7. MonkeyTree Prototype
Prototype microbenchmark.   Figure 14 presents the timing breakdown of migrating a single worker in a $4$-node Llama3-8B training instance using our prototype. The top column plots overhead on the sender side (source of migration), while the bottom one plots the receiver side (destination of migration). The dominant cost is composing the checkpoint at $3.93$ s. MonkeyTree-daemon transfers the $12$ GB checkpoint shard (model weights and optimizer state) over RDMA in $3.19$ s end-to-end. The transfer itself accounts for only $1.07$ s at $90$ Gbps; the remainder is serialization and memory management overhead. Source server cleanup adds $1.24$ s, though this does not block the migrated model from resuming. On the destination, NCCL initialization, model loading, and other startup costs total $2.97$ s. Excluding the RDMA transfer’s network time, the total overhead is $9.02$ s, validating the $10$ s migration budget excluding RDMA transfer assumed in simulations.
@@ 8. Discussion
Extending to three-tier topologies.   The partitioning strategy above enables a natural extension to three-tier (or deeper) topologies: the second-tier MonkeyTree controller manages fragmentation beyond the spine switches. Furthermore, an extension of Theorem 1 shows that if no single job exceeds the bounds of a partition, then $1$ is achievable for cross-partition fragmentation. Therefore, only one uplink per partition is necessary. This keeps the routing problem polynomial, enabling a straightforward extension of Perfect routing to three-tier topologies. We leave the performance evaluation of this scheme to future work.

### v2

@@ 2.1. Multi-Tenant Training and Congestion
$2.$ Migration Challenges in CPU Workloads
@@ 3. Tractability of GPU Migration
DP replicates the model across GPUs, with each GPU processing a different data batch. Fully Sharded Data Parallelism (FSDP) extends DP by sharding parameters across GPUs. Both require synchronizing gradients or parameters at the end of each training iteration through collective communication. These collectives generate traffic proportional to model size with the same volume and pattern, so we treat them identically and refer to both as DP traffic. When combined with other parallelisms such as $P$ and TP, only a subset of GPUs shares the same model replica; we call this subset a replica group, and each performs its own collective independently.
@@ 3.1. Congestion-Free States are Achievable
This sparsity resolves the ill-defined optimization problem from Section $2.$ In particular, there exist states in which the traffic leaving each ToR does not exceed its uplink capacity (bottom-left of Figure 3). We define such states as congestion-free, since this condition allows each DP flow to be assigned a dedicated uplink, achieving path isolation. Under minimal uplink provisioning, such states always exist (§5.1). For CPU workloads, operators must balance network contention against migration overhead (joint_migration_routing; joint_migration_routing_2; joint_placement_routing_3). In contrast, for ML workloads this tradeoff disappears: because congestion-free states are always attainable, they can be enforced as a constraint, allowing the optimization to focus solely on minimizing migrations.
@@ 3.2. Congestion-Free States Are Abundant
$3.$ Congestion-Free States are Self-Reinforcing
@@ 5.2. MonkeyTree ILP for DP and FSDP Jobs
We attribute this efficiency to the ability of branch-and-bound solvers (branch_bound) to exploit the abundance of the solution space. To build intuition for the runtime, let $\gamma$ denote the number of migrations in searching for a solution. The number of distinct cluster states reachable within $\gamma$ moves is $O(N^{2\gamma})$, where NN is the cluster size. As branch-and-bound progressively discovers feasible solutions with increasing migration counts, large portions of the search space are pruned, since the solver discards solutions requiring the same or more moves. We believe that, because the optimal number of migrations is typically small (§6.5), the solver explores a limited effective search space, resulting in fast ILP solving time.
@@ 5.3. Extending to Model Parallel Jobs
Pipeline parallelism ($P$).   Pipeline parallelism shards the model between layers into distinct stages that run on separate GPUs. Activations are communicated point-to-point between stages. Data-parallel replicas are thus across each stage, and the subsequent all-reduce (DP-flows) only happens within these stages. We only use our Perfect routing algorithm to load balance pipeline traffic and ignore accommodating them in our formulation, because they are small (§3) and therefore have little impact on end-to-end job completion time. Instead, we update our formulation to treat pipeline stages as independent jobs, thereby controlling each stage’s DP-flow separately.
@@ 6.5. Migration Solver Scalability
Figure 9 presents the distribution of required job migration (number of moves) to lower the fragmentation below two on clusters of varying sizes across five traces each. The distribution shifts as cluster size increases, but the average remains consistently between $1.7$ and $1.8$. Approximately $80\%$ of the weight is concentrated on one or two moves across all scenarios, and move counts greater than 55 are rare, occurring in fewer than $1\%$ of the migrations for each cluster.
@@
Finally, Figure 12 plots the ILP’s solve time across cluster sizes ranging from $128$ to $2,048$ GPUs. For clusters of $1,024$ GPUs or fewer, the average solve time remains under $0.5$ s with a p90 below 11 s. At $2,048$ GPUs, solve time increases by over $5\times$, with an average of $2.39$ s and a p90 of $4.04$ s. Even so, this remains well within practical limits: migrations are triggered infrequently and can tolerate seconds of planning latency, making the ILP viable even at the large cluster sizes.
@@ 6.7. Threshold Sensitivity
Figure 13 plots the number of migrations alongside the average, p90, and p99 slowdown. At a threshold of 44, matching the number of uplinks, p99 slowdown is within 66% of ideal, and the average is within 11%. This is slightly worse than the results in §6.2, as a workload composed entirely of FSDP and DP jobs with few uplinks presents a more constrained setting than the full workload mix. Decreasing the threshold below four yields less than 0.50.5% additional improvement: Perfect routing already guarantees path isolation at a fragmentation threshold of four, so further migrations cannot eliminate elephant-flow collisions. A lower threshold only reduces collisions between elephant flows and in-flight migration traffic, yielding a marginal benefit offset at the cost of up to $2\times$ more migrations.
@@ 7. MonkeyTree Prototype
Prototype microbenchmark.   Figure 14 presents the timing breakdown of migrating a single worker in a 44-node Llama3-8B training instance using our prototype. The top column plots overhead on the sender side (source of migration), while the bottom one plots the receiver side (destination of migration). The dominant cost is composing the checkpoint at $3.93$ s. MonkeyTree-daemon transfers the $12$ GB checkpoint shard (model weights and optimizer state) over RDMA in $3.19$ s end-to-end. The transfer itself accounts for only $1.07$ s at $90$ Gbps; the remainder is serialization and memory management overhead. Source server cleanup adds $1.24$ s, though this does not block the migrated model from resuming. On the destination, NCCL initialization, model loading, and other startup costs total $2.97$ s. Excluding the RDMA transfer’s network time, the total overhead is $9.02$ s, validating the $10$ s migration budget excluding RDMA transfer assumed in simulations.
@@ 8. Discussion
Extending to three-tier topologies.   The partitioning strategy above enables a natural extension to three-tier (or deeper) topologies: the second-tier MonkeyTree controller manages fragmentation beyond the spine switches. Furthermore, an extension of Theorem 1 shows that if no single job exceeds the bounds of a partition, then 11 is achievable for cross-partition fragmentation. Therefore, only one uplink per partition is necessary. This keeps the routing problem polynomial, enabling a straightforward extension of Perfect routing to three-tier topologies. We leave the performance evaluation of this scheme to future work.